OpenStudy (anonymous):

Fill in the reason that justifies the step by step. Given: QS = 42 Prove: x = 13

OpenStudy (samwisegamge):

............ummmmmmmm

OpenStudy (anonymous):

QS= 42 QR+RS= QS (x+3) + 2x=42 3x+3=42 3x=39 x=13

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

can you help?